OpenStudy (anonymous):

The ordered pairs (1,81), (2,100), (3,121),(4,144), and (5,169)represent a function. What is a rule that represents this function? (1 point)

OpenStudy (alekos):

81 = 9^2, 100 = 10^2, 121 = 11^2 and so on from here you can work out a relationship between the pairs

OpenStudy (dayakar):

81 = 9^2, 100 = 10^2, 121 = 11^2 ,144 =12^2, 169 = 13^2, \[f \left( x \right)= \left( x+8 \right)^{2}, if x=1,2,3,4,5\]